# BTL-UI: Blink-Think-Link Reasoning Model for GUI Agent

### v2

@@ 4.1 Implementation Details
Experimental Setup. We develop the BTL-UI-3B/7B model based on Qwen2.5-VL-3B/7B and adopt the ms-swift zhao2024swiftascalablelightweightinfrastructure framework for RL training. As shown in Table 1, we train BTL-UI on a grounding and planning mixture. All experiments are conducted on NVIDIA H100-80G GPUs.
@@
Grounding ShowUI-Desktop lin2024showui 1K
@@
Grounding: Screenspot series benchmarks assess fundamental GUI understanding and element grounding accuracy across diverse platforms (Mobile, Desktop, Web). ScreenSpot cheng2024seeclick evaluates single-step GUI grounding performance across multiple platforms. ScreenSpot-V2 cheng2024seeclick , a re-annotated version, addresses annotation errors present in the original ScreenSpot.
Table 2: GUI grounding accuracy on ScreenSpot cheng2024seeclick . Bold means the best results, and underline means the second best results.
@@
Evaluation Metrics. For grounding tasks, we use click point prediction accuracy as our evaluation metric. For planning tasks, in line with OS-Atlas wu2024atlas , we report three standard metrics for GUI agents: action type prediction accuracy (Type), click point prediction accuracy (GR), and step success rate (SR). Specifically: Type measures the exact‐match accuracy between predicted and ground‑truth action types (e.g., “click” vs. “swipe”). GR evaluates grounding performance via click point prediction accuracy in specific action types (e.g., “click” and “long press”). SR is the step‑wise success rate: a step is counted as successful only if both the predicted action and its associated arguments (e.g., click coordinates or input text) match the ground truth.
@@ 4.2 Experimental Results
We evaluate BTL-UI across three key capabilities: grounding, low-level planning, and high-level reasoning. The results demonstrate consistent and significant improvements over existing baselines, validating the effectiveness of the Blink-Think-Link framework. For more experimental results, please refer to the supplemental materials.
Table 3: GUI grounding accuracy on ScreenSpot-V2 cheng2024seeclick . Bold means the best results, and underline means the second best results.
@@
Grounding Capability. To assess how well the model can localize UI elements, we report grounding accuracy on the ScreenSpot benchmark series in Table 2 and 3. On the original ScreenSpot dataset, BTL-UI-7B achieves an average accuracy of 87.2%, outperforming the baseline Qwen2.5-VL-7B (84.8%), and surpassing the supervised fine-tuned Aria-UI (82.4%). On the corrected ScreenSpot-V2, performance further improves to 89.1%, establishing a new SOTA. This suggests that the Blink Phase, which encourages early-stage attention to semantically relevant regions through ROI supervision, enables more precise perception and grounding even under diverse visual layouts. Notably, even the 3B variant achieves highly competitive results, indicating strong sample efficiency of our training paradigm.
Low-Level Planning Capability. On AndroidControl (low-level), which requires accurate step-by-step GUI interactions, BTL-UI-3B reaches a step success rate (SR) of 84.8%. This outperforms both GUI-R1-3B (64.4%), and the best SFT model—SeeClick (75.0%). Moreover, BTL-UI-7B achieves SOTA in all metrics of AndroidControl-Low. BTL-UI reduces common errors such as off-target clicks or premature task termination. The agent shows robustness to UI variations and fine-grained action transitions, a core challenge in real-world control scenarios.
High-Level Planning Capability. For long-horizon tasks on AndroidControl (high-level) and GUI-Odyssey, BTL-UI-7B achieves 69.2% SR and 45.2% SR, respectively, outperforming GUI-R1-7B by large margins (51.7% and 35.2%). These tasks require the agent not only to ground visual entities but also to reason through abstract goals and perform multi-step sequences. The Blink Phase plays a pivotal role by generating targeted ROI annotations at each decision point, effectively filtering out visual clutter and guiding the policy toward relevant elements early in the planning process. The Think Phase contributes here by providing reasoning scaffolds and structured action decomposition, while the Link Phase ensures execution fidelity.
@@ 4.3 Ablation Study
As shown in Table 5, to clarify the contributions of each component in our BTL framework, we conduct an ablation study on the AndroidControl-High benchmark. When trained only with Supervised Fine-Tuning (SFT), BTL-UI achieves a baseline performance with an SR of 60.6%. While further using the generated Blikn data, SFT obtains a 5% improvement. This proves that BLink data is not only suitable for RFT, but also for SFT. Furthermore, RFT without BLink data archives an SR of 65.6%. After adopting Blink data and BTL reward, SR is improved by 3.8%.
@@
Overall, the ablation results confirm that each element of the BTL framework—Blink Phase for targeted attention, Think Phase for structured reasoning, and the Link Phase for precise validation—plays a crucial role in achieving state-of-the-art performance in GUI interaction tasks.
@@ 4.4 Visualization
Due to space limitations, we present the visualization results and qualitative analysis in the supplementary material.
@@ 5 Conclusion and Limitations
We propose the BTL framework, an innovative GUI interaction architecture inspired by the biological cognitive paradigm of "Blink-Think-Link". This framework simulates the human closed-loop system of visual perception, cognitive decision-making, and action execution during GUI operations, overcoming the limitations of traditional outcome-driven RFT approaches. Experimental results show that the BTL-UI agent, developed under this framework, achieves significant performance improvements across a variety of GUI interaction tasks.
We believe that the BTL framework proposed in this study establishes a novel paradigm for developing digital assistants that are more natural, efficient, and aligned with human cognition. It not only benefits human-GUI interaction but can also be extended to other human-computer interaction tasks, such as embodied intelligence.
Limitations. The proposed BTL framework introduces <blink> tag outputs compared to conventional Think-Answer structured outputs. Although the blink-generated ROI regions are adaptive and can be empty (zero-length), they typically increase the output sequence length in most cases. While demonstrating performance improvements across various GUI task metrics, this design incurs additional processing overhead.
Appendix A Addition Experimental Results
Table 6: GUI grounding accuracy on Screenspot-Pro [41]. Bold means the best results, and underline means the second best results.
Model Method Model Size Dev Creative CAD Scientific Office OS Avg.
Text Icon Text Icon Text Icon Text Icon Text Icon Text Icon
GPT-4o [38] ZS - 1.3 0.0 1.0 0.0 2.0 0.0 2.1 0.0 1.1 0.0 0.0 0.0 0.8
Qwen2-VL [4] ZS 7B 0.5 0.0 2.6 0.0 1.5 0.0 6.3 0.0 3.4 1.9 0.9 0.0 1.6
Qwen2.5-VL [5] ZS 3B - - - - - - - - - - - - 23.9
Qwen2.5-VL [5] ZS 7B - - - - - - - - - - - - 29.0
OS-Atlas-Base [23] ZS 7B 33.1 1.4 28.8 2.8 12.2 4.7 37.5 7.3 33.9 5.7 27.1 4.5 18.9
ShowUI [35] SFT 2B 16.9 1.4 9.1 0.0 2.5 0.0 13.2 7.3 15.3 7.5 10.3 2.2 7.7
UGround [22] SFT 7B 26.6 2.1 27.3 2.8 14.2 1.6 31.9 2.7 31.6 11.3 17.8 0.0 16.5
UGround-V1 [22] SFT 7B - - - - - - - - - - - - 31.1
SeeClick [7] SFT 9.6B 0.6 0.0 1.0 0.0 2.5 0.0 3.5 0.0 1.1 0.0 2.8 0.0 1.1
CogAgent [6] SFT 18B 14.9 0.7 9.6 0.0 7.1 3.1 22.2 1.8 13.0 0.0 5.6 0.0 7.7
UI-R1 [30] RFT 3B 11.2 6.3 22.7 4.1 27.3 3.5 42.4 11.8 32.2 11.3 13.1 4.5 17.8
GUI-R1 [10] RFT 3B 26.4 7.8 33.8 4.8 40.9 5.6 61.8 17.3 53.6 17.0 28.1 5.6 -
GUI-R1 [10] RFT 7B 23.9 6.3 49.4 4.8 38.9 8.4 55.6 11.8 58.7 26.4 42.1 16.9 -
BTL-UI RFT 7B 53.9 7.3 54.4 15.9 35.9 14.6 47.2 13.0 62.7 24.7 55.7 19.7 33.7
To further evaluate the grounding capability of BTL-UI, we evaluate it on Screenspot-Pro [41].
Screenspot-Pro focuses on high-resolution professional environments, featuring expert-annotated tasks spanning 23 applications, five industries, and three operating systems. The experimental result on Screenspot-Pro is shown in Table 6. Our proposed BTL‑UI models achieve the best performance across both 3B and 7B model scales. Notably, the 7B version of BTL‑UI consistently outperforms all other models, achieving an average accuracy of 33.7%, which is substantially higher than GUI-R1 (23.9%) and UI-R1 (17.8%) at the same scale. This highlights the effectiveness of our task-aligned adaptation and grounding-oriented training based on Qwen2.5-VL. In particular, BTL‑UI‑7B exhibits strong performance in user-intensive scenarios such as Dev, Office, and OS, achieving icon grounding accuracies of 7.3%, 24.7%, and 19.7%, respectively—far exceeding existing baselines.
Appendix B Prompt
@@
The prompt of BTL-UI is shown in Table 7. The system prompt is used to format the output of the model according to the three-phase paradigm of blink-think-link. Moreover, the model outputs according to the format of the system prompt, which is convenient for the calculation of the BTL reward to adjust the model distribution. As shown in equation 6, due to the output of Blink Phase can be $\varnothing$, we emphasize that the Blink Phase can be output None in the system prompt.
OS-Atlas [23] has found that blindly mixing data from different sources for multitask fine-tuning can significantly harm performance due to action space conflicts. We unify the action space of GUI-Odyssey [37] and AndroidControl [36]. And we prompt the model to select the corresponding action from the defined action space. Furthermore, we declare the format and purpose in the user instruction prompt, so that the model can better understand each action type. For the grounding and high-level tasks, only the high-level instruction will be fed into the model. While for the low-level tasks, both the high-level instructions and the low-level instructions will be given to the model. In the end, the interaction history will be added to the user instruction prompt.
Appendix C Visualization
@@
The interaction trajectory visualization of our BTL-UI is shown in Figure 4. The high-level instruction is “Listen live to Radio GupShup 94.3 FM and search for other radio stations.” The Blink Phase can locate the ROIs related to the instruction. And the Think Phase can reason based on the instruction, interaction history, and the candidate area. As shown in step 2 of the interaction trajectory in Figure 4, in the Blink Phase, BTL-UI not only locates the input box to complete the task, but also analyzes the historical search records in the screenshots.
@@
Appendix D Experiment Statistical Significance

### v3

@@ 4.1 Implementation Details
Experimental Setup. We develop the BTL-UI-3B/7B model based on Qwen2.5-VL-3B/7B and adopt the ms-swift framework zhao2024swiftascalablelightweightinfrastructure for RL training. As shown in Table 1, we train BTL-UI in a mix of grounding and planning data.
@@
Grounding ShowUI-Web lin2024showui 1K
@@
Grounding: Screenspot series benchmarks assess fundamental GUI understanding and element grounding accuracy across diverse platforms (Mobile, Desktop, Web). ScreenSpot cheng2024seeclick evaluates the single-step GUI grounding performance across multiple platforms. ScreenSpot-V2 cheng2024seeclick , a re-annotated version, addresses annotation errors present in the original ScreenSpot. ScreenSpot-Pro li2025screenspot specifically increases the difficulty with complex desktop applications and high-resolution screens.
Table 2: GUI grounding accuracy on ScreenSpot cheng2024seeclick . Bold means the best results, and underline means the second best results. Avg. denotes the average performance on mobile, desktop and web subtasks.
@@
UGround-V1 gou2024navigating SFT 2B 89.4 72.0 88.7 65.7 81.3 68.9 77.7
UGround-V1 gou2024navigating SFT 7B 94.1 79.9 93.8 76.4 90.9 84.0 86.3
UI-TARS qin2025ui SFT 2B 93.0 75.5 90.7 68.6 84.3 74.8 82.3
UI-TARS qin2025ui SFT 7B 94.5 85.2 95.9 85.7 90.0 83.5 89.5
@@
Evaluation Metrics. For grounding tasks, we use click point prediction accuracy as our evaluation metric. For planning tasks, according to OS-Atlas wu2024atlas , we report three standard metrics for GUI agents: action type prediction accuracy (Type), click point prediction accuracy (GR) and step success rate (SR). Specifically: Type measures the exact‐match accuracy between predicted and ground‑truth action types (e.g., “click” vs. “swipe”). GR evaluates grounding performance via click point prediction accuracy in specific action types (e.g. “click” and “long press”). SR is the step‑wise success rate: a step is counted as successful only if both the predicted action and its associated arguments (e.g., click coordinates or input text) match the ground truth.
@@ 4.2 Experimental Results
We evaluate BTL-UI across three key capabilities: grounding, low-level planning, and high-level reasoning. The results demonstrate consistent and significant improvements over existing baselines, validating the effectiveness of the Blink-Think-Link framework.
Table 3: GUI grounding accuracy on ScreenSpot-V2 cheng2024seeclick . Bold means the best results, and underline means the second best results. Avg. denotes the average performance on mobile, desktop and web subtasks.
@@
Grounding Capability. To assess how well the model can localize UI elements, we report grounding accuracy on the ScreenSpot benchmark series in Table 2, 3, and 4. In the original ScreenSpot dataset, BTL-UI-7B achieves an average accuracy of 87.2%, outperforming the baseline Qwen2.5-VL-7B (84.8%) and surpassing the supervised fine-tuned Aria-UI (82.4%). On the corrected ScreenSpot-V2, performance further improves to 89.1%. In the ScreenSpot-Pro benchmark, the BTL‑UI-3B consistently outperforms other RFT-based models, achieving an average accuracy of 27.1%, which is substantially higher than UI-R1 (17.8%) on the same scale. This suggests that the Blink Phase, which encourages early-stage attention to semantically relevant regions through ROI supervision, enables more precise perception and grounding even under diverse visual layouts. Although the overall grounding performance of BTL-UI remains slightly lower than that of UI-TARS qin2025ui , which is a strong GUI Agent developed based on Qwen2-VL wang2024qwen2 with training on 50B tokens, the proposed BTL-UI shows certain advantages in the mobile subtasks.
Planning Capability. As shown in Table 5, BTL-UI exhibits strong generalization and reasoning ability on both low-level and high-level GUI planning benchmarks. In AndroidControl-Low, BTL-UI-7B achieves an SR of 88.0%, surpassing the previous best model OS-Atlas-Pro-7B (85.2%) and GUI-R1-7B (66.5%) , while the 3B variant attains a comparable 84.8%, confirming the efficiency of the BTL reinforcement paradigm. For long-horizon tasks in AndroidControl-High, which require multi-step reasoning and contextual grounding, BTL-UI-7B achieves an SR of 69.2%, outperforming GUI-R1-7B (51.7%) and approaching the SFT-based GUI foundation model, OS-Atlas-Pro-7B (71.2%). This improvement reflects the synergy between Blink-phase attention and Link-phase symbolic reward, which jointly stabilize execution and reduce accumulated errors in extended interaction sequences. In GUI-Odyssey, a benchmark that emphasizes hierarchical decision-making and interface switching, BTL-UI-7B reaches an SR of 65.4%, significantly surpassing GUI-R1-7B (38.8%). Although there is still a performance gap compared to UI-TARS, the proposed BTL-UI is comparable to large-scale SFT models such as OS-Atlas, while requiring significantly less training data.
Table 4: GUI grounding accuracy on Screenspot-Pro li2025screenspot . Bold means the best results, and underline means the second best results. Avg. denotes the average performance on all subtasks.
Model Method Model Size Dev Creative CAD Scientific Office OS Avg.
Text Icon Text Icon Text Icon Text Icon Text Icon Text Icon
GPT-4o openai2024hello ZS - 1.3 0.0 1.0 0.0 2.0 0.0 2.1 0.0 1.1 0.0 0.0 0.0 0.8
Qwen2-VL wang2024qwen2 ZS 7B 0.5 0.0 2.6 0.0 1.5 0.0 6.3 0.0 3.4 1.9 0.9 0.0 1.6
Qwen2.5-VL bai2025qwen2 ZS 3B - - - - - - - - - - - - 23.9
Qwen2.5-VL bai2025qwen2 ZS 7B - - - - - - - - - - - - 29.0
OS-Atlas-Base wu2024atlas ZS 7B 33.1 1.4 28.8 2.8 12.2 4.7 37.5 7.3 33.9 5.7 27.1 4.5 18.9
ShowUI lin2024showui SFT 2B 16.9 1.4 9.1 0.0 2.5 0.0 13.2 7.3 15.3 7.5 10.3 2.2 7.7
UGround gou2024navigating SFT 7B 26.6 2.1 27.3 2.8 14.2 1.6 31.9 2.7 31.6 11.3 17.8 0.0 16.5
UGround-V1 gou2024navigating SFT 7B - - - - - - - - - - - - 31.1
SeeClick cheng2024seeclick SFT 9.6B 0.6 0.0 1.0 0.0 2.5 0.0 3.5 0.0 1.1 0.0 2.8 0.0 1.1
CogAgent hong2024cogagent SFT 18B 14.9 0.7 9.6 0.0 7.1 3.1 22.2 1.8 13.0 0.0 5.6 0.0 7.7
UI-TARS qin2025ui SFT 2B 47.4 4.1 42.9 6.3 17.8 4.7 56.9 17.3 50.3 17.0 21.5 5.6 27.7
UI-TARS qin2025ui SFT 7B 58.4 12.4 50.0 9.1 20.8 9.4 63.9 31.8 63.3 20.8 30.8 16.9 35.7
UI-R1 lu2025ui RFT 3B 11.2 6.3 22.7 4.1 27.3 3.5 42.4 11.8 32.2 11.3 13.1 4.5 17.8
GUI-R1 xia2025gui RFT 3B 26.4 7.8 33.8 4.8 40.9 5.6 61.8 17.3 53.6 17.0 28.1 5.6 -
GUI-R1 xia2025gui RFT 7B 23.9 6.3 49.4 4.8 38.9 8.4 55.6 11.8 58.7 26.4 42.1 16.9 -
BTL-UI RFT 7B 53.9 7.3 26.7 15.9 35.9 14.6 47.2 13.0 62.7 24.7 55.7 19.7 33.7
@@
OS-Atlas-Pro wu2024atlas SFT 4B 91.9 83.8 80.6 84.7 73.8 67.5 83.5 61.4 56.4
OS-Atlas-Pro wu2024atlas SFT 7B 93.6 88.0 85.2 85.2 78.5 71.2 84.5 67.8 62.0
UI-TARS qin2025ui SFT 2B 98.1 87.3 89.3 81.2 78.4 68.9 93.9 86.8 83.4
UI-TARS qin2025ui SFT 7B 98.0 89.3 90.8 83.7 80.5 72.5 94.6 90.1 87.0
@@ 4.3 Ablation Study
As shown in Table 6, to clarify the contributions of each component in our BTL framework, we conduct an ablation study on the AndroidControl-High benchmark. When trained only with SFT, BTL-UI achieves a baseline performance with an SR of 60.6%. While further using the generated Blink data, SFT obtains a 5% improvement. This proves that Blink data is not only suitable for RFT, but also for SFT. Furthermore, RFT without Blink data achieves an SR of 65.6%. After adopting Blink data and BTL reward, SR is improved by 3.6%.
@@
Overall, the ablation results confirm that each element of the BTL framework—Blink Phase for targeted attention, Think Phase for structured reasoning, and the Link Phase for precise validation—plays a crucial role in achieving competitive performance in GUI interaction tasks.
@@ 4.4 Visualization
We present the visualization results and qualitative analysis in the appendix.
@@ 5 Conclusion and Limitations
We propose the BTL framework, an innovative GUI interaction architecture inspired by the biological cognitive paradigm of Blink–Think–Link. This framework simulates the human closed-loop system of visual perception, cognitive decision-making, and action execution during GUI operations, overcoming the limitations of traditional outcome-driven RFT approaches. Experimental results show that the BTL-UI agent, developed under this framework, achieves significant performance improvements across a variety of GUI interaction tasks.
We believe that the BTL framework proposed in this study establishes a promising and generalizable paradigm for developing digital assistants that are more natural, efficient, and aligned with human cognition. It not only benefits human-GUI interaction but can also be extended to other human-computer interaction tasks, such as embodied intelligence.
Limitations. The proposed BTL framework introduces <blink> tag outputs compared to conventional Think-Answer structured outputs. Although the blink-generated ROI regions are adaptive and can be empty (zero-length), they typically increase the output sequence length in most cases. While demonstrating performance improvements across various GUI task metrics, this design incurs additional computational processing overhead.
Appendix A Prompt
@@
The prompt of BTL-UI is shown in Table 7. The system prompt is used to format the output of the model according to the three-phase paradigm of Blink-Think-Link. Moreover, the model outputs according to the format of the system prompt, which is convenient for the calculation of the BTL reward to adjust the model distribution. As shown in equation 6, because the output of Blink Phase can be $\varnothing$, we emphasize that the Blink Phase can be output None in the system prompt.
OS-Atlas [23] has found that blindly mixing data from different sources for multitask fine-tuning can significantly harm performance due to action space conflicts. We unify the action space of GUI-Odyssey [37] and AndroidControl [36]. And we prompt the model to select the corresponding action from the defined action space. Furthermore, we declare the format and purpose in the user instruction prompt, so that the model can better understand each action type. For the grounding and high-level tasks, only the high-level instruction will be fed into the model. For the low-level tasks, both the high-level and the low-level instructions will be given to the model. In the end, the interaction history will be added to the user instruction prompt.
Appendix B Visualization
@@
The visualization of the interaction trajectory of our BTL-UI is shown in Figure 4. The high-level instruction is “Listen live to Radio GupShup 94.3 FM and search for other radio stations.” The Blink Phase can locate the ROIs related to the instruction. And the thinking Phase can reason based on the instruction, interaction history, and candidate area. As shown in step 2 of the interaction trajectory in Figure 4, in the Blink Phase, BTL-UI not only locates the input box to complete the task, but also analyzes the historical search records in the screenshots.
@@
Appendix C Experiment Statistical Significance